What is the difference between probably and perhaps?
An action modified by perhaps is less likely than the action modified by probably.
1.It Meant Music Performed For Dancing The Polka...
2.a A Bohemian Dance With 3 Steps And a Hop In Fast Time...
If two dice are rolled simultaneously what is the probability of getting a 1 and a 2?
The total number of combinations are 6 x 6 or 36. Each of these combinations is equally likely. You can make a table with six rows and six columns to see all combinations. Two combinations, (1,2) and (2,1) will satisfy getting a one on one die and a two on the other one. Therefore the probability = 2/36 or 1/18.
probability of hitting a white key 7/12, probability of hitting a black key 5/12. Assuming all key have an equal possibility of being hit.
In a bag of 40 marbles 18 are blue what percent of marbles in the bag are blue?
make it out of 100, so 40 times 2.5 equals 100 and 18 times 2.5 equals... 45 so its 45 percent
What is the probability of drawing a five or a jack?
The probability of drawing a five or a jack from a standard deck of 52 cards is 8 in 52, or 2 in 13, or about 0.1538.
What is the probability a boy is chosen if their are 26 kids and 10 are girls?
If 10 out of 26 are girls, then the probability of randomly choosing a boy is 16 out of 26, or 8 out of 13, or about 0.6154.
Is it possible to roll a pair of six sided dice and get fourteen?
A six-sided die only goes up to 6. Therefore, when rolling two, you can get a total of 2 times 6, so 12 is the highest number possible, and you cannot roll a fourteen.
This is a combination problem because you have to assess a subgroup without a particular order. I find it easiest to first deduce what the total number of arrangements are and then subtract from there.
Firstly, There are a total of 9 balls getting assigned to a group of 4. So, to get our total, we use 9C4, which equals 9!/(4!*5!)=126. This means there are a total of 126 arrangements regardless of order. However, this number also includes the combination of 4 like-colored balls we're told to avoid.
To calculate the number of arrangements for 4 like-colored balls we set up a separate subgroup for each color. For red balls the possible arrangement are 5C4, which is 5!/(4!*1!)=5. For blue balls the possible arrangements are 4C4, which is 4!/(4!*0!)=1 (remember that 0!=1). So the total number of like-colored arrangements is 5+1=6.
Now we just subtract the exceptions from the total to get the answer...126-6=120
In the standard normal distribution the means is always?
0
This may not make sense, but it takes a concept known as z-scores into account, which basically standardizes everything.
What events occurred on December 7th?
Japanese bomb Pearl Harbor, Hawaii; also attack the Philippines, Wake Island, Guam, Malaya, Thailand, Shanghai and Midway.
if repeating is allowed than 1/263= 1/17576, if repeating isn't allowed than it's 1/(26*25*24)=1/15600.
A coin is tossed two times what is the probability of obtaining at least one tail?
3/4
The only way a tail doesn't show up is if all heads come up.
P(all heads) = (1/2)*(1/2)=1/4
P(at least one tail)=1-P(all heads)=1-(1/4)=3/4
Assuming you mean from a standard 52-card deck of four suits, there is a 4/52 chance of first dealing an ace, a 4/51 chance of secondly dealing a king and a 4/50 chance of dealing a queen. Together, these mean there is a 64/132600, or 8/16575 or 0.04826546% chance of dealing the desired cards in order.
How are identical twins genetically from fraternal twins?
Identical twins are genetically the same, because they come from the same fertilized egg/sperm cell zygote, which happens to split into two viable zygotes.
Fraternal twins, however, come from two different egg/sperm cell zygotes, that are concurrently fertilized and remain viable.
What are the odds of rolling a straight 1-6 using 5 dice in one roll?
Impossible, you can't get 1-6 with only 5 dice.
Rolling three dice what is the probability of rolling an odd number?
The probability that, when rolling three dice none will come up odd is broken down by the probability that on odd roll will occur as an independent event. Rolling the first die has a probability of 3/6 that it will be odd. Rolling the second has the probability of 3/6 that it will be odd. Rolling the third has a probability of 3/6 will be odd. Rolling three and coming up with odd is really the probability that no odd numbers occurred, so therefore an even number DID happen (3/6), and an even number DID happen (3/6), and an even number DID happen (3/6). That's really just 1 - (3/6^3) or simplified 1 - (1/2^3). So therefore the probability that an odd number occurred in 3 die rolls is about 87.5%.
What does probability mean in math terms?
you are finding the probable answer in a question
EXAMPLE: Michelle has 10 cards. She is giving 2 to Shannon. What would the fraction, the dot (decimal) and the percent be?
ANSWER: The fraction is... 2-10
The decimal is... .20
The percent is...20%
What is the theoretical probability of throwing a coin and it landing on heads?
Theoretical probability is the number of ways something can occur divided by the total number of outcomes. So, the theoretical probability of throwing a coin and it landing on heads is 1/2 or 0.5 or 50%.
Consider a coin toss. The probability of the coin coming up head is 1/2, the probability of the coin coming up tails is 1/2. No matter how many times you flip the coin the probability of any particular toss coming up heads or tails is always 1/2.
BUT
If you consider two coins tosses as a set then you have four possible outcomes:
first coin heads - second coin heads, first coin heads - second coin tails,
first coin tails - second coin heads, first coin tails - second coin tails
for ease of demonstration lets denote heads as H and tails as T
the above combination would then be HH,HT,TH,TT
the probability of any one of the combinations would be 1/2 for the first coin and 1/2 for the second coin so all combinations have a probability of 1/4. (1/2x1/2=1/4 I hope you realize.)
But, if we don't care which coin is H or T then the two combinations HT and TH are the same and we combined their probabilities into one 1/4+1/4=1/2
HH,(HT,TH),TT
1/4 1/2 1/4
Two dice rolled what is the probability of getting 16?
Assuming a 6 sided die, a roll of 2 gives a maximum value of 12. So the probability of getting a 16 is zero.
Example of statistics problem using binomial theorem?
The classic one is the coin toss problem. How many head and how many tails on a certain number of tosses.
For example we toss a coin 4 times and call a success the result of the coin landing with the head showing. The binomial variable is the number of heads n the four coin tosses, which can take on the values 0, 1, 2, 3, or 4. The probabilities of the possible outcome can be calculated using the binomial theorem.
Here they are:
P(X = 0) = 1/16
P(X = 1) = 1/4
P(X = 2) = 3/8
P(X = 3) = 1/4
P(X = 4) = 1/16
P = 0.5n (n!/(n-x)!x! )
What does replacement mean in probability?
When the sample is drawn, it is placed back where it was taken from and if subsequent draws are made, it could be selected again.